OpenStudy (joshoyen):

Write a polynomial function of least degree with integral coefficients that has the given zeros. 1. -3 + √10, -3 - √10, √5

OpenStudy (danjs):

degree 4 with Integer ?

OpenStudy (danjs):

you remember the zeros of a polynomial are the solutions of y = P(x) = 0

OpenStudy (danjs):

and when you factored sometimes, you had a product of quantities , if any one was 0, then y=0 ( x + )(x + ).....=0

OpenStudy (joshoyen):

oh so is it (+3-√10) ?

OpenStudy (danjs):

this is the reverse, make quantities using each of those given zeroes, \[(x - 1)*(x - \sqrt{5}) = 0\] that takes care of 2 of the 4,

OpenStudy (danjs):

when x=1, you get p(x)=0 , also when x is root5,

OpenStudy (danjs):

if those are complex number pair, is there an imaginary i in them , or they just real numbers like that

OpenStudy (danjs):

\[(x - 1)*(x - \sqrt{5})*(x + (-3 + \sqrt{10})*(x + (-3 - \sqrt{10})) = 0\] that there works for all them zero values given for x. you just need to expand all that out into standard polynomial , most likely

OpenStudy (joshoyen):

ohhh okay, i see now thanks!

OpenStudy (danjs):

welcome